OpenStudy (anonymous):

What are the vertices of the ellipse given by the equation the quantity x plus 3 end quantity squared over 16 plus the quantity y minus 4 end quantity squared over 121= 1?

OpenStudy (anonymous):

(–3, 8) and (–3, 0) (–3, –7) and (–3, 15) (–14, 4) and (8, 4) (1, 4) and (–7, 4)

OpenStudy (jdoe0001):

so \(\bf \cfrac{(x-h)^2}{a^2}+\cfrac{(y-k)^2}{b^2}=1\\ \cfrac{(x+3)^2}{16}+\cfrac{(y-4)^2}{121} = 1 \implies \cfrac{(x+3)^2}{4^2}+\cfrac{(y-4)^2}{11^2} = 1\) what do you think is the (h, k) point for the center of the ellipse?

OpenStudy (anonymous):

h=3 and k=4 right?

OpenStudy (jdoe0001):

well, keep in mind that (x+3) => ( x - ( - 3) )

OpenStudy (jdoe0001):

and (y - 4) => ( y - ( +4 ) )

OpenStudy (anonymous):

so h= 3 and k= -4?

OpenStudy (jdoe0001):

the standard equation for an ellipse is \(\bf \cfrac{(x-h)^2}{a^2}+\cfrac{(y-k)^2}{b^2}=1\)

OpenStudy (anonymous):

right so where do i go from there

OpenStudy (jdoe0001):

well, what's your center?

OpenStudy (jdoe0001):

\(\bf \cfrac{(x+3)^2}{4^2}+\cfrac{(y-4)^2}{11^2} = 1 \implies \cfrac{(x-(-3))^2}{4^2}+\cfrac{(y-(+4))^2}{11^2} = 1\\ \cfrac{(x-\color{red}{h})^2}{a^2}+\cfrac{(y-\color{red}{k})^2}{b^2}=1\)

OpenStudy (anonymous):

3,-4?

OpenStudy (jdoe0001):

notice the 2nd version

OpenStudy (anonymous):

-3, 4? sorry i dont follow

OpenStudy (jdoe0001):

the standard equation is (x-h) if you have (x + 3) and the "x" component would be 3, only if the standard equation would use (x +h), but it doesn't, it uses (x-h) that means that for (x+3) to turn into a (x - h) form, it has to be ( x - ( -h) ) minus times minus = plus

OpenStudy (jdoe0001):

so the center is at (-3, 4) = ( h, k) the "a" component of an ellipse is the bigger SQUARED denominator and it's also the major axis, and the vertices will be THAT MUCH FAR FROM THE center of the ellipse whichever variable that has that bigger denominator, is where the ellipse is LONGER at in your case, "y"

OpenStudy (jdoe0001):